OpenStudy (anonymous):

Simplify the expression. [9 + (–5)] • (–4) A. 16 B. –56 C. 56 D. –16

OpenStudy (anonymous):

9-5x-4=-16

OpenStudy (imstuck):

There's no x in there, though.

OpenStudy (anonymous):

-16

OpenStudy (imstuck):

9 + (-5) = 9-5 = 4. 4 * -4 = -16

OpenStudy (anonymous):

Wht do you mean? I wanted to check if it was -16

OpenStudy (anonymous):

It is -16